OpenStudy (anonymous):

A radio station has a broadcast area in the shape of a circle with equation x^2 + y^2 = 5,625, where the constant represents square miles.

hartnn (hartnn):

whats the question ?

OpenStudy (anonymous):

A radio station has a broadcast area in the shape of a circle with equation x^2 + y^2 = 5,625, where the constant represents square miles.

OpenStudy (anonymous):

I'm very confused on it.

hartnn (hartnn):

thats just an information, anything to fin out ??

hartnn (hartnn):

*find

hartnn (hartnn):

please post entire Question...

OpenStudy (anonymous):

a. Graph the equation and state the radius in miles. b. What is the area of the region in which the broadcast from the station can be picked up?

hartnn (hartnn):

x^2+y^2 =r^2 is the circle with centre = (0,0) and radius = r

hartnn (hartnn):

so, x^2 + y^2 = 5,625 is the circle with centre = (0,0) and radius \(= \sqrt{5625}=...?\)

OpenStudy (anonymous):

75?

hartnn (hartnn):

yes, thats the radius in miles. r=75 miles. then area = \(\pi r^2\)

OpenStudy (anonymous):

17662.5?

OpenStudy (anonymous):

is that right?

hartnn (hartnn):

accurately, its 17671.46 sq. miles.

OpenStudy (anonymous):

Is that the answer?

hartnn (hartnn):

yup, for part b)

OpenStudy (anonymous):

I figured out par A already. Could you help me with a few more, every time I ask my teacher for help she gets frustrated with me?

OpenStudy (anonymous):

part*

hartnn (hartnn):

can i request you to ask in separate post ? so that if i am unavailable, others can help you....

OpenStudy (anonymous):

Yes I just posted another